OpenStudy (anonymous):

A bag contains 20 checkers --10 red and 10 black. Find the probability of selecting 2 black checkers when they are chosen at random without replacement. Determine whether the event is dependent or independent.

OpenStudy (kropot72):

The probability of selecting a black checker on the first choice is 10/20. Having selected a black checker on the first choice, there are 9 black checkers remaining in a total of 19 checkers. So the probability of selecting a black checker on the second choice is 9/19. The probability of selecting 2 black checkers is the product of the above two values of probability.

OpenStudy (anonymous):

so what would the answer be?

OpenStudy (kropot72):

\[P(2\ black\ checkers)=\frac{10}{20}\times\frac{9}{19}=you\ can\ calculate\]

OpenStudy (anonymous):

i got 9/38 but i am confused on if it would be dependent or independent

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

it is dependent, because the probability the second one is black depends on whether or not the first one was black

OpenStudy (anonymous):

i am guessing of course this line Determine whether the event is dependent or independent. makes no sense at all one event is not "dependent" or "independent" two events are, it is a comparison

OpenStudy (anonymous):

either it is a typo, or the person who wrote the question is confused about probability

OpenStudy (anonymous):

are you pretty positive in your guess?

OpenStudy (anonymous):

i think its asking if the answer is independent or dependent